OpenStudy (anonymous):

Use basic identities to simplify the expression. cosine of theta to the second power divided by sine of theta to the second power. + csc θ sin θ

OpenStudy (anonymous):

@Elsa213 I don't understand

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Hero @IMStuck

zepdrix (zepdrix):

Simba again!! :O What's the monkey's name? I don't remember +_+

OpenStudy (anonymous):

Rafikki lol any idea how to do this one? I have a list of identities but idk which to use

zepdrix (zepdrix):

\(\Large\rm \frac{\cos^2\theta}{\sin^2\theta}+\csc\theta \sin\theta\) this is the problem?

OpenStudy (anonymous):

yep

zepdrix (zepdrix):

Remember your cosecant identity? \(\Large\rm \csc\theta=\frac{1}{\sin\theta}\) That should help us simplify the second term.

zepdrix (zepdrix):

Rafikki? Mm ok yah that sounds right :d Scar was the cool one. so scury

OpenStudy (anonymous):

could we make that part sin^2 by multiplying 1/sin and sin by sin?

OpenStudy (anonymous):

I'm a personal fan of Kovu (scar's son in Lion King 2) :p

zepdrix (zepdrix):

Oh ive never seen that one :o

zepdrix (zepdrix):

Our second term we want to plug in our identity for csc(theta) \(\Large\rm \color{orangered}{\csc\theta}\sin\theta=\color{orangered}{\frac{1}{\sin\theta}}\sin\theta\) Multiplying, \(\Large\rm \frac{\sin\theta}{\sin\theta}\) and cancel stuff out, you end up with 1 for your second term, yes?

zepdrix (zepdrix):

For the first term, write it like this: \(\Large\rm \left(\frac{\cos\theta}{\sin\theta}\right)^2\) then recall your identity, (cosx/sinx)=cotx

OpenStudy (anonymous):

so we're left with just cotx?

zepdrix (zepdrix):

Yes, but with a square on it, \(\Large\rm \cot^2\theta+1\) And we were left with a 1 from the second term

zepdrix (zepdrix):

From there you can apply your Pythagorean Identity that involves cotangent and cosecant.

OpenStudy (anonymous):

oh and cot^2+1=csc^2!

zepdrix (zepdrix):

yay gj \c:/

OpenStudy (anonymous):

got time for one more similar one?

zepdrix (zepdrix):

sorry class ending :c i gotta pack up

OpenStudy (anonymous):

aww okay, thanks for the help!